A roulette wheel has 38 slots: 18 red, 18 black, and 2 green. The ball ends at one of these at random.

You are a player who plays a large number of games and makes an even bet of Ghc 1 on red in every

game. After n games, what is the probability that you are ahead? Answer this for n = 100 and

n = 1000.



1
Expert's answer
2021-05-07T14:39:34-0400

Let "S_n" be the number of times you win. This is a Binomal "(n,\dfrac{9}{19})" random variable.


P(ahead)=P(win more than half of the games)


     "=P(S_n>\dfrac{n}{2})
\\[9pt]
 =P(\dfrac{S_n-np}{\sqrt{np(1-p)}}>\dfrac{\frac{n}{2}-np}{\sqrt{np(1-p)}}
)
\\[9pt] =P(Z>\dfrac{(\frac{1}{2}-p)\sqrt{n}}{\sqrt{p(1-p)}})"

  

For n=100,

"P(Z>\dfrac{5}{\sqrt{90}}=P(Z>0.527)=0.2990"


For n=1000.

"P(Z>\dfrac{5}{3})=P(Z>1.66)=0.0478"